Find the value(s) of y such that the triangle with the given vertices has an area of 8 square units. (Enter your answers as a comma-separated list.)
(−4, 5), (−3, 4), (−1, y).

Answer by Alan3354(69443)

Find the value(s) of y such that the triangle with the given vertices has an area of 8 square units. (Enter your answers as a comma-separated list.)
A(−4,5), B(−3,4), C(−1,y).
-----------------

 A    B    C   A
-4   -3   -1  -4
 5    4    y   5

Add the diagonal products starting at the upper left.
-4*4 + -3y -1*5 = -3y - 21
---
Add the diagonal products starting at the lower left.
5*-3 + 4*-1 - 4y = -4y - 19
---
The ABS of the difference is y - 2
The area is 1/2 that = y/2 - 1
----
y/2 - 1 = 8
y = 18
==================
The other difference is -y + 2
-y/2 + 1 = 8
y = -14